How to graph a linear system and estimate the solution. Then check the solution algebrically.
problem: 2x + y= 13
5x- 2y= 1

Answer by elima(1433)

2x + y= 13
5x- 2y= 1
There are different was to solve this. We will solve it by elimination;
First multiplt the eq.1 by 2;
(2)2x+y=13
4x+2y=26
Now we add the eq.1 and eq.2 together so we can eliminate the y.
4x+2y=26
5x-2y=1
----------
9x=27
x=3
Now we now x=3, so we can plug that into the equation and get y.
2(3)+y=13
6+y=13
y=7
Now go back and check your answer;
2(3)+(7)=13
6+7=13
It checks out!!!
Hope you understand
=)